# Learning to Ask: When LLMs Meet Unclear Instruction

### v2

@@ Abstract
Equipped with the capability to call functions, modern large language models (LLMs) can leverage external tools for addressing a range of tasks unattainable through language skills alone. However, the effective execution of these tools relies heavily not just on the advanced capabilities of LLMs but also on precise user instructions, which often cannot be ensured in the real world. To evaluate the performance of LLMs tool-use under imperfect instructions, we meticulously examine the real-world instructions queried from users, analyze the error patterns, and build a challenging tool-use benchmark called Noisy ToolBench (NoisyToolBench). We find that due to the next-token prediction training objective, LLMs tend to arbitrarily generate the missed argument, which may lead to hallucinations and risks. To address this issue, we propose a novel framework, Ask-when-Needed (AwN), which prompts LLMs to ask questions to users whenever they encounter obstacles due to unclear instructions. Moreover, to reduce the manual labor involved in user-LLM interaction and assess LLMs’ performance in tool utilization from both accuracy and efficiency perspectives, we design an automated evaluation tool named ToolEvaluator. Our experiments demonstrate that the AwN significantly outperforms existing frameworks for tool learning in the NoisyToolBench. We will release all related code and datasets to support future research.
Introduction
@@
LLMs have undergone remarkable development since OpenAI introduced ChatGPT-3.5 (Bang et al. 2023). This model demonstrates a significant advancement in solving multiple tasks, including code generation (Dong et al. 2023; Sakib, Khan, and Karim 2023; Feng et al. 2023), machine translation (Jiao et al. 2023; Peng et al. 2023), even game playing (Wu et al. 2024). However, despite their impressive capabilities, LLMs often struggle with complex computations and delivering accurate, timely information (Qu et al. 2024). Tool learning emerges as a promising solution to mitigate these limitations of LLMs by enabling dynamic interaction with external tools (Schick et al. 2024).
@@
The incorporation of tool usage capabilities marks a pivotal step towards enhancing the intelligence of LLMs, pushing them closer to exhibiting human-like intelligence. The integration of tool usage allows LLMs to perform a broader array of complex and varied tasks, including managing emails, designing presentations, and browsing the web to gather real-time information. For example, LLMs can perform complex calculations using a calculator tool, access real-time weather updates through weather APIs, and execute programming code via interpreters (Qin et al. 2023a; Schick et al. 2024; Mialon et al. 2023; Yang et al. 2023a).
Toolformer (Schick et al. 2024) is a pioneering work in empowering language models with self-learning capabilities for tool usage. This groundbreaking approach integrates various tools, including calculators, Q&A systems, and search engines. Then, significant research efforts have been directed toward accessing a wider variety of tools or using multiple tools simultaneously to resolve a single query, such as Gorilla(Patil et al. 2023), RestGPT (Song et al. 2023) and ToolLLM (Qin et al. 2023b).
@@
Figure 1: Motivating Example of Our Method.
Despite the significant strides made, existing frameworks and benchmarks often operate under the assumption that user instructions are always explicit and unambiguous, a premise that diverges from real-world scenarios (Qin et al. 2023a; Song et al. 2023; Patil et al. 2023). Due to the feature of API calls, it requires precise user instructions since the arguments for the function call can hardly tolerate ambiguity. We find that due to the next-token prediction training objective, LLMs tend to arbitrarily generate the missed argument, which may lead to hallucinations and risks (as the example shown in Figure 1a). Furthermore, as the tasks assigned to LLMs grow in complexity, multiple and sequential API calls are needed to resolve a single task. This complexity amplifies the challenge, as any error in the sequence of API calls can culminate in an outcome that strays from the user’s original intention. Hence, LLMs tool-use under unclear instruction is an important but rarely investigated direction.
To address this oversight, we conduct a systematic analysis of actual user instructions, identifying and categorizing potential issues into several key areas. These include instructions lacking essential information, instructions with ambiguous references, instructions containing inaccuracies, and instructions that are unfeasible for LLMs to execute due to the limitations of the tools available. Building on this observation, we meticulously design a noisy instruction benchmark for tool use, named NoisyToolBench. This benchmark includes a collection of provided APIs, ambiguous queries, anticipated questions for clarification, and the corresponding responses. Its primary goal is to assess the capability of LLMs to detect ambiguities in user queries and to pose relevant questions for clarification accordingly.
@@
To evaluate the performance of LLMs tool-use under unclear instruction, we design several innovative metrics from the accuracy and efficiency perspectives. For accuracy, we measure the LLMs’ proficiency in asking appropriate clarifying questions, their ability to execute the correct function calls, and their success in delivering final responses that meet the users’ needs. For efficiency, we calculate the average number of redundant asked questions and the average number of actions required to complete the instruction. An ideal LLM should achieve higher accuracy with fewer number of queries. Recognizing the labour-intensive nature of manually communicating with LLMs and verifying all execution results, we also innovatively design an automatic evaluation system, ToolEvaluator, to streamline the whole process. ToolEvaluator leverages the advanced problem-solving capabilities of GPT-4 to communicate with LLMs and automatically evaluate the performance of LLMs’ tool-using under unclear instruction. Our experiments demonstrate that the AwN significantly outperforms existing baseline methods for tool learning in the NoisyToolBench.
@@
  • We conduct a systematic examination of the real-world user instruction for tool utilization and categorize the prevalent issues into four distinct categories.
@@
  • We design five evaluation metrics from both accuracy and efficiency perspectives and introduce an automatic evaluation system, ToolEvaluator, that can proxy humans to interact with and assess LLMs.
@@
  • We introduce a novel algorithm, named AwN method, to prompt LLMs to actively ask questions to request clarifications from users when facing uncertainties. Experimental results show that AwN can significantly improve the LLMs’ tool-using under unclear instructions.
Related Works
@@
Tool Learning for LLMs. LLMs have recently made significant advancements, with ChatGPT being recognized as a major step towards achieving AGI (Wu et al. 2023; Lund and Wang 2023; Jiao et al. 2023). These LLMs possess strong reasoning capabilities, enabling them to perform increasingly complex tasks (Liu et al. 2023). However, to progress further towards AGI, it is crucial for LLMs to master the utilization of tools. Toolformer is the first innovative AI model designed to use several specialized tools, such as a web browser, a code interpreter, and a language translator, within a single framework (Schick et al. 2023). The model’s ability to seamlessly switch between these tools and apply them contextually represents a significant advancement in AI capabilities. Recent studies like RestGPT (Song et al. 2023) and ToolLLM (Qin et al. 2023b), have connected LLMs with real-life Application Programming Interfaces (APIs), allowing them to sequentially employ multiple external tools to solve user queries. The tool-augmented approach empowers LLMs to use various kinds of tools to do more sophisticated tasks, showcasing an enhanced level of capability compared to pure LLMs. Besides, API-Bank (Li et al. 2023), ToolAlpaca (Tang et al. 2023), ToolBench (Qin et al. 2023b), ToolQA (Zhuang et al. 2023) and RestBench (Song et al. 2023) are exemplary benchmarks to systematically evaluate the performance of tool-augmented LLMs performance in response to user’s queries. However, current models often ignore the situations in which users might not give exact instructions, which can result in the tools not working properly. Thus, our study aims to tackle this specific challenge by developing a new benchmark specifically for ambiguous instructions.
@@
Prompting LLMs for Decision Making. In certain situations, addressing user queries may require more than a single API call. This necessitates the effective division of the overarching task into smaller, more manageable components, which presents a significant challenge. Prior research has focused extensively on enhancing LLMs’s ability to effectively plan and execute complex tasks. The ’Chain of Thought’ prompting approach facilitates advanced reasoning by introducing intermediate steps in the reasoning process (Wei et al. 2022). The ReAct methodology improves the integration of reasoning and action, enabling LLMs to take informed actions based on environmental feedback (Yao et al. 2022). Meanwhile, Reflexion is designed to reduce errors in the reasoning process by revisiting and learning from previous mistakes (Shinn, Labash, and Gopinath 2023). DFSDT expands upon Reflexion, allowing LLMs to evaluate various options and choose the most viable path (Qin et al. 2023b). In our work, we innovatively involve users in the process of executing instructions. Our approach, referred to as AwN, motivates LLMs to consider the necessity of requesting further information from users during each tool invocation round. This strategy aims at clarifying users’ ambiguous instructions to help execute the tasks in alignment with the users’ intentions.
@@
Learning to Ask. Since user queries may not always be clear, and the execution of LLMs may encounter uncertainties and ambiguities, learning to ask questions has emerged as a challenging yet crucial research area. Some researchers introduce a learning framework that empowers an embodied visual navigation agent to proactively seek assistance(Zhang et al. 2023). Recently, similar ideas have been adopted in the software engineering, leveraging a communicator to enhance the reliability and quality of generated code (Wu 2023). Our work focuses on the tool-learning scenario, which is more sensitive to the user’s unclear query. A concurrent study (Qian et al. 2024) also focuses on the reliability of tool-learning systems under unclear instruction. However, they did not systematically examine real-world user behavior, leading to the limited and biased nature of their dataset that doesn’t accurately capture user errors. Our research addresses this shortfall by starting with a user analysis. Additionally, Qian’s methodology depends significantly on human manual interaction and assessment of LLM performances, which is time-consuming and hard to reproduce. In contrast, we introduce an automated evaluation method that can proxy humans to communicate with and automatically evaluate the performance of LLMs.
Noisy ToolBench
@@
Several tool-learning benchmarks have been introduced to assess LLMs’ ability in tool utilization. However, these benchmarks overlook the potential ambiguity in users’ instruction, which might hinder LLMs from executing tasks as intended by the user. For instance, as depicted in Figure 1a, if a user inquires, ”How is today’s weather” without specifying the location, LLMs cannot accurately activate the APIs to fetch the correct weather information. This scenario underscores the critical role of interaction between users and LLMs in executing instructions accurately. However, previous tool-learning benchmarks only contain perfect user instruction in a one-query-one-execution manner.
@@
User Instruction Analysis
To analyze the issues in real-world user instruction, we recruit human annotators to write user queries according to the API provided. Firstly, we select 100 APIs from the ToolBench (Qin et al. 2023b), containing real-world RESTful APIs spanning 49 categories, ranging from sports to finance. Secondly, we hire 10 volunteers, who have a Bachelor’s degree, are proficient in English, and have experience using LLMs. We provide them with the 100 APIs, and then ask them to write down an instruction to prompt LLMs to call each API, ending up with 1000 user queries. Finally, we manually identify the problematic user queries and categorized them as follows.
  • Instructions Missing Key Information (IMKI): These are user instructions that omit crucial details necessary for the successful execution of a function. An example of IMKI would be, ”Set an alarm to wake me up” without providing a specific time. Asking for more information is needed when encountering this issue.
  • Instructions with Multiple References (IMR): These user instructions include elements that can be interpreted in several ways, potentially leading to confusion for LLMs in understanding the user’s actual intent. For example, an IMR instance is ”I want to know the director of the movie ’The Matrix’,” where the ambiguity arises because there are multiple versions of ’The Matrix’, each possibly having a different director. This issue is similar to IMKI but is more subtle and difficult to detect. Pointing out potential references and asking for clarification are needed when encountering this issue.
  • Instructions with Errors (IwE): This category consists of user instructions that contain the necessary information for executing a function, but the information is incorrect. An example of IWE is, ”Please help me to log in to my Twitter. My user account is ’abcde@gmail.com’ and the password is ’123456’,” where the user might have provided the wrong account details or password due to typographical errors. Asking for the correct information is needed when encountering this issue.
@@
Table 1 shows the ratio of the four issues, where the most common issue in the instructions is ”Instructions Missing Key Information”, with a significant 56% of all errors. This issue is a clear indication that users often do not provide adequate information to effectively use the APIs. Additionally, issues such as ”Instructions with Errors” and ”Instructions Beyond Tool Capabilities” were identified at rates of 17.3% and 15.3%, respectively.
@@
Instructions Missing Key Information (IMKI) 56%
@@
Table 1: The distribution of various problematic instructions.
Data Construction
Our analysis of user instructions reveals that there are four kinds of instruction issues that may lead to LLMs’ tool utilization failures: Instructions Missing Key Information (IMKI), Instructions with Multiple References (IMR), Instructions with Errors (IwE), and Instructions Beyond Tool Capabilities (IBTC). Consequently, we build our benchmark with the four issues by intentionally modifying the problem-free instructions from well-established datasets to problematic ones. We first select 200 data with problem-free instruction from ToolBench and then manually modify the user instructions to make them suffer from the four kinds of instruction issues. Then we annotate the expected questions that LLMs should ask when facing each imperfect user query, which will be used to measure whether LLMs can ask the right questions, as well as the answer to the question, which will be used to proxy the human responses. We conduct a two-round cross-verification to ensure the quality of the annotation. Each data is annotated and verified by different people and any disagreement data will be re-annotated until reach a consensus. Finally, each data entry in NoisyToolBench has the following five components: the imperfect user query, the available APIs, the questions that LLMs should ideally ask, the answers to these questions, and the expected function calls along with their respective arguments.
Figure 2: A comparison of our QwN prompting compared with original CoT/ReAct Prompting
Ask-when-Needed Prompting

### v3

# Learning to Ask: When LLM Agents Meet Unclear Instruction

@@ Abstract
Equipped with the capability to call functions, modern LLM agents can leverage external tools for addressing a range of tasks unattainable through language skills alone. However, the effective execution of these tools relies heavily not just on the advanced capabilities of LLM agents but also on precise user instructions, which often cannot be ensured in the real world. To evaluate the performance of LLM agents tool-use under imperfect instructions, we meticulously examine the real-world instructions queried from users, analyze the error patterns, and build a challenging tool-use benchmark called Noisy ToolBench (NoisyToolBench). We find that due to the next-token prediction training objective, LLM agents tend to arbitrarily generate the missed argument, which may lead to hallucinations and risks. To address this issue, we propose a novel framework, Ask-when-Needed (AwN), which prompts LLM agents to ask questions to users whenever they encounter obstacles due to unclear instructions. Moreover, to reduce the manual labor involved in user-LLM interaction and assess LLM agents’ performance in tool utilization from both accuracy and efficiency perspectives, we design an automated evaluation tool named ToolEvaluator. Our experiments demonstrate that the AwN significantly outperforms existing frameworks for tool learning in the NoisyToolBench. We will release all related code and datasets to support future research.
Learning to Ask: When LLM Agents Meet Unclear Instruction
Wenxuan Wang1  Juluan Shi1thanks:   Denote Equal Contribution.  Zixuan Ling1  Yuk-Kit Chan1  Chaozheng Wang1 Cheryl Lee1  Youliang Yuan2  Jen-tse Huang 1thanks:   Jen-tse Huang is the corresponding author.  Wenxiang Jiao2  Michael R. Lyu1 1The Chinese University of Hong Kong   2Tencent AI Lab 3The Chinese University of Hong Kong, Shenzhen 1{wxwang,jthuang,lyu}@cse.cuhk.edu.hk2joelwxjiao@tencent.com
1 Introduction
@@
LLMs have undergone remarkable development since OpenAI introduced ChatGPT-3.5 Bang et al. (2023). This model demonstrates a significant advancement in solving multiple tasks, including code generation Dong et al. (2023); Sakib et al. (2023); Feng et al. (2023), machine translation Jiao et al. (2023); Peng et al. (2023), even game playing Wu et al. (2024). However, despite their impressive capabilities, LLMs often struggle with complex computations and delivering accurate, timely information Qu et al. (2024). Tool learning emerges as a promising solution to mitigate these limitations of LLMs by enabling dynamic interaction with external tools Schick et al. (2024).
@@
The incorporation of tool usage capabilities marks a pivotal step towards enhancing the intelligence of LLMs, pushing them closer to exhibiting human-like intelligence. The integration of tool usage allows LLMs to perform a broader array of complex and varied tasks, including managing emails, designing presentations, and browsing the web to gather real-time information. For example, LLMs can perform complex calculations using a calculator tool, access real-time weather updates through weather APIs, and execute programming code via interpreters Qin et al. (2023a); Schick et al. (2024); Mialon et al. (2023); Yang et al. (2023a). Toolformer Schick et al. (2024) is a pioneering work in empowering language models with self-learning capabilities for tool usage. Then, significant research efforts have been directed toward accessing a wider variety of tools or using multiple tools simultaneously to resolve a single query, such as GorillaPatil et al. (2023), RestGPT Song et al. (2023) and ToolLLM Qin et al. (2023b).
@@
Figure 1: The motivating example of our Ask-when-Needed (AwN) framework.
Despite the significant strides made, existing frameworks and benchmarks often operate under the assumption that user instructions are always explicit and unambiguous, a premise that diverges from real-world scenarios Qin et al. (2023a); Song et al. (2023); Patil et al. (2023). Due to the feature of API calls, it requires precise user instructions since the arguments for the function call can hardly tolerate ambiguity. We find that due to the next-token prediction training objective, LLMs tend to arbitrarily generate the missed argument, which may lead to hallucinations and risks (as the example shown in Figure 1a). Furthermore, as the tasks assigned to LLMs grow in complexity, multiple and sequential API calls are needed to resolve a single task. This complexity amplifies the challenge, as any error in the sequence of API calls can culminate in an outcome that strays from the user’s original intention. Hence, LLMs tool-use under unclear instruction is an important but rarely investigated direction.
To address this oversight, we conduct a systematic analysis of actual user instructions, identifying and categorizing potential issues into several key areas. These include instructions lacking essential information, instructions with ambiguous references, instructions containing inaccuracies, and instructions that are unfeasible for LLMs to execute due to the limitations of the tools available. Building on this observation, we meticulously design a noisy instruction benchmark, named NoisyToolBench, which is pimarily used for assessing the capability of LLMs to detect ambiguities in user queries and to pose relevant questions for clarification accordingly. Specifically, NoisyToolBench includes a collection of provided APIs, ambiguous queries, anticipated questions for clarification, and the corresponding responses.
@@
To evaluate the performance of LLMs tool-use under unclear instruction, we design several innovative metrics from the accuracy and efficiency perspectives. For accuracy, we measure the LLMs’ proficiency in asking appropriate clarifying questions, their ability to execute the correct function calls, and their success in delivering final responses that meet the users’ needs. For efficiency, we calculate the average number of redundant asked questions and the average number of actions required to complete the instruction. An ideal LLM should achieve higher accuracy with fewer number of queries. Recognizing the labour-intensive nature of manually communicating with LLMs and verifying all execution results, we also innovatively design an automatic evaluation system, ToolEvaluator, to streamline the whole process. ToolEvaluator leverages the advanced problem-solving capabilities of GPT-4o to communicate with LLMs and automatically evaluate the performance of LLMs’ tool-using under unclear instruction. Our experiments on 6 LLMs and 2 tool-using frameworks demonstrate that the AwN significantly outperforms existing baseline methods for tool learning in the NoisyToolBench.
@@
  • We conduct a systematic study on real-world user instruction for tool utilization and categorize the prevalent issues into four distinct categories.
@@
  • We design five evaluation metrics from both accuracy and efficiency perspectives and introduce an automatic evaluation system, ToolEvaluator, that can proxy users to interact and assess LLMs.
@@
  • We introduce a novel framework, named AwN method, to prompt LLMs to actively ask questions to request clarifications from users when facing uncertainties. Experimental results show that AwN can significantly improve the LLMs’ tool-using under unclear instructions.
2 Related Works
@@
Tool Learning for LLMs. LLMs have recently made significant advancements, with ChatGPT being recognized as a major step towards achieving AGI Wu et al. (2023); Lund and Wang (2023); Jiao et al. (2023). These LLMs possess strong reasoning capabilities, enabling them to perform increasingly complex tasks Liu et al. (2023). However, to progress further towards AGI, it is crucial for LLMs to master the utilization of tools. Toolformer is the first innovative AI model designed to use several specialized tools, such as a web browser, a code interpreter, and a language translator, within a single framework Schick et al. (2023). The model’s ability to seamlessly switch between these tools and apply them contextually represents a significant advancement in AI capabilities. Recent studies like RestGPT Song et al. (2023) and ToolLLM Qin et al. (2023b), have connected LLMs with real-life Application Programming Interfaces (APIs), allowing them to sequentially employ multiple external tools to solve user queries. The tool-augmented approach empowers LLMs to use various kinds of tools to do more sophisticated tasks, showcasing an enhanced level of capability compared to pure LLMs. Besides, API-Bank Li et al. (2023), ToolAlpaca Tang et al. (2023), ToolBench Qin et al. (2023b), ToolQA Zhuang et al. (2023) and RestBench Song et al. (2023) are exemplary benchmarks to systematically evaluate the performance of tool-augmented LLMs performance in response to user’s queries. However, current models often ignore the situations in which users might not give exact instructions, which can result in the tools not working properly. Thus, our study aims to tackle this specific challenge by developing a new benchmark specifically for ambiguous instructions.
@@
Prompting LLMs for Decision Making. In certain situations, addressing user queries may require more than a single API call. This necessitates the effective division of the overarching task into smaller, more manageable components, which presents a significant challenge. Prior research has focused extensively on enhancing LLMs’s ability to effectively plan and execute complex tasks. The ’Chain of Thought’ prompting approach facilitates advanced reasoning by introducing intermediate steps in the reasoning process Wei et al. (2022). The ReAct methodology improves the integration of reasoning and action, enabling LLMs to take informed actions based on environmental feedback Yao et al. (2022). Meanwhile, Reflexion is designed to reduce errors in the reasoning process by revisiting and learning from previous mistakes Shinn et al. (2023). DFSDT expands upon Reflexion, allowing LLMs to evaluate various options and choose the most viable path Qin et al. (2023b). In our work, we innovatively involve users in the process of executing instructions. Our approach, referred to as AwN, motivates LLMs to consider the necessity of requesting further information from users during each tool invocation round. This strategy aims at clarifying users’ ambiguous instructions to help execute the tasks in alignment with the users’ intentions.
@@
Learning to Ask. Since user queries may not always be clear, and the execution of LLMs may encounter uncertainties and ambiguities, learning to ask questions has emerged as a challenging yet crucial research area Rao and Daumé III (2018); Kuhn et al. (2022); Andukuri et al. (2024). For example, some researchers introduce a learning framework that empowers an embodied visual navigation agent to proactively seek assistanceZhang et al. (2023). Recently, similar ideas have been adopted in the software engineering, leveraging a communicator to enhance the reliability and quality of generated code Wu (2023). Our work focuses on the tool-learning scenario, which is more sensitive to the user’s unclear query. A concurrent study Qian et al. (2024) also focuses on the reliability of tool-learning systems under unclear instruction. However, they did not systematically examine real-world user behavior, leading to the limited and biased nature of their dataset that doesn’t accurately capture user errors. Our research addresses this shortfall by starting with a user analysis. Additionally, Qian’s methodology depends significantly on human manual interaction and assessment of LLM performances, which is time-consuming and hard to reproduce. In contrast, we introduce an automated evaluation method that can proxy humans to communicate with and automatically evaluate the performance of LLMs.
3 Noisy ToolBench
@@
Several tool-learning benchmarks have been introduced to assess LLMs’ ability in tool utilization. However, these benchmarks overlook the potential ambiguity in users’ instruction, which might hinder LLMs from executing tasks as intended by the user. For instance, as depicted in Figure 1a, if a user inquires, "How is today’s weather" without specifying the location, LLMs cannot accurately activate the APIs to fetch the correct weather information. This scenario underscores the critical role of interaction between users and LLMs in executing instructions accurately. However, previous tool-learning benchmarks only contain perfect user instruction in a one-query-one-execution manner.
@@
3.1 User Instruction Analysis
To analyze the issues in real-world user instruction, we recruit human annotators to write user queries according to the API provided. Firstly, we select 100 APIs from the ToolBench Qin et al. (2023b), containing real-world RESTful APIs spanning 49 categories, ranging from sports to finance. Secondly, we hire 10 volunteers, who have a Bachelor’s degree, are proficient in English, and have experience using LLMs. We provide them with the 100 APIs, and then ask them to write down an instruction to prompt LLMs to call each API, ending up with 1000 user queries. Finally, we manually identify the problematic user queries and categorized them as follows.
  • Instructions Missing Key Information (IMKI): These are user instructions that omit crucial details necessary for the successful execution of a function. An example of IMKI would be, "Set an alarm to wake me up" without providing a specific time. Asking for more information is needed when encountering this issue.
  • Instructions with Multiple References (IMR): These user instructions include elements that can be interpreted in several ways, potentially leading to confusion for LLMs in understanding the user’s actual intent. For example, an IMR instance is "I want to know the director of the movie ’The Matrix’," where the ambiguity arises because there are multiple versions of ’The Matrix’, each possibly having a different director. This issue is similar to IMKI but is more subtle and difficult to detect. Pointing out potential references and asking for clarification are needed when encountering this issue.
  • Instructions with Errors (IwE): This category consists of user instructions that contain the necessary information for executing a function, but the information is incorrect. An example of IWE is, "Please help me to log in to my Twitter. My user account is ’abcde@gmail.com’ and the password is ’123456’," where the user might have provided the wrong account details or password due to typographical errors. Asking for the correct information is needed when encountering this issue.
@@ 3.1 User Instruction Analysis
Table 1 shows the ratio of the four issues, where the most common issue in the instructions is "Instructions Missing Key Information", with a significant 56.0% of all errors. This issue is a clear indication that users often do not provide adequate information to effectively use the APIs. Additionally, issues such as "Instructions with Errors" and "Instructions Beyond Tool Capabilities" were identified at rates of 17.3% and 15.3%, respectively.
@@
Instructions Missing Key Information (IMKI) 56.0%
@@
Table 1: Distribution of problematic instructions.
3.2 Data Construction
Our user instruction analysis reveals that there are four kinds of instruction issues that may lead to LLMs’ tool utilization failures: Instructions Missing Key Information (IMKI), Instructions with Multiple References (IMR), Instructions with Errors (IwE), and Instructions Beyond Tool Capabilities (IBTC). So, we build our benchmark with the four issues by intentionally modifying the problem-free instructions from well-established datasets to problematic ones. We first select 200 data with problem-free instruction from ToolBench and then manually modify the user instructions to make them suffer from the four kinds of instruction issues. Then we annotate the expected questions that LLMs should ask when facing each imperfect user query, which will be used to measure whether LLMs can ask the right questions, as well as the answer to the question, which will be used to proxy the human responses. We conduct a two-round cross-verification to ensure the quality of the annotation. Each data is annotated and verified by different people and any disagreement data will be re-annotated until reach a consensus. Finally, each data entry in NoisyToolBench has the following five components: the imperfect user query, the available APIs, the questions that LLMs should ideally ask, the answers to these questions, and the expected function calls along with their respective arguments.
Figure 2: The comparison of our QwN prompting compared with original CoT/ReAct Prompting
4 Ask-when-Needed Prompting